Express cash flow as uniform payment series for 10
years that is equivalent to a payment series of Rs. 50,000 at the end of first year, decreasing by Rs.
5,000 each year over 10 years at interest rate of 16% per year compounded annually.
1
Expert's answer
2020-05-28T11:07:55-0400

50000-5000=45 000


Postnumerando Annuity


"FVA=А\times\frac{(1+r)^n-1}{r}=45 000\times\frac{(1+0.16)^{10}-1}{0.16}=959466.11"